OpenStudy (anonymous):

During the 1998-1999 Little League season, the Tigers played 52 games. They won 12 more games than they lost. How many games did they win that season?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

2x-12=52 2x=52+12 2x=64 x=32

OpenStudy (dumbcow):

w+L = 52 w = L+12 substitute L+12 for w in 1st equation (L+12)+L = 52 2L +12 = 52 2L = 40 L = 20 w = 20+12 = 32

OpenStudy (anonymous):

hoped this helped

OpenStudy (anonymous):

Hey dumbcow do you mind helping me with mine

OpenStudy (anonymous):

Thank you all for responding, I've been sitting for a while trying to figure this out(:

OpenStudy (dumbcow):

just did

OpenStudy (anonymous):

well your in luck i had this similar problem 2 wks ago on a test

OpenStudy (anonymous):

Thank you DumbCow

OpenStudy (anonymous):

This is one of my online college math class and I'm not all that math savy :P